OpenStudy (anonymous):

x^2+x-2/(x-1)(x+3) determine the vertical asymtotes

OpenStudy (shubhamsrg):

set denominator = 0 , find respective values of x. Those values of will represent your vertical asymptotes.

OpenStudy (anonymous):

As vertical asymptotes are the y-coordinates reaching positive or negative infinity, we would expect \[y\rightarrow \pm \infty\] so \[\frac{x^2+x-2}{(x-1)(x+3)}\rightarrow \pm \infty\] hence \[\frac{1}{(x-1)(x+3)}\rightarrow\pm\infty\] This implies \[(x-1)(x+3)=0\] Then solve for the value for x and you would have \[x=1\]or\[x=-3\]